# Simple is Effective: The Roles of Graphs and Large Language Models in Knowledge-Graph-Based Retrieval-Augmented Generation

### v2

@@ 4.1 Evaluation for Retrieval (Q1 & Q2)
WebQSP-sub CWQ-sub Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh G-Retriever 54.13 23.84 74.52 67.56 67.97 - - - - - RoG-Joint 72.01 47.70 88.90 82.62 76.13 58.61 52.12 66.22 61.17 55.15 RoG-Sep 67.94 43.10 84.03 77.61 72.79 57.69 52.83 64.64 60.64 54.51 SubgraphRAG + Llama3.1-8B 72.10 46.56 88.58 84.80 82.42 54.76 51.76 65.80 59.69 62.89 SubgraphRAG + Llama3.1-70B 75.97 51.64 87.88 85.89 85.57 61.49 59.91 68.43 65.52 67.62 SubgraphRAG + ChatGPT 70.81 44.73 85.18 80.82 81.53 56.37 54.44 64.40 60.99 61.31 SubgraphRAG + GPT-4o-mini 78.34 58.44 91.34 87.36 82.21 61.13 58.86 70.01 65.48 64.20 SubgraphRAG + GPT-4o 77.61 56.78 91.40 86.40 81.85 65.99 63.18 73.91 68.89 66.57 SubgraphRAG + GPT-4o-mini (200) 78.66 58.65 91.73 87.04 81.98 61.58 57.47 71.45 65.87 63.66 SubgraphRAG + GPT-4o (200) 79.40 58.91 92.43 87.75 82.46 66.48 61.30 75.14 69.42 66.45 SubgraphRAG + GPT-4o-mini (500) 78.46 57.08 92.43 88.01 81.95 62.18 56.86 72.82 66.57 62.77 SubgraphRAG + Llama3.1-8B ($\leftrightarrow$) 67.91 42.79 85.25 81.21 80.09 43.03 40.73 55.09 47.58 56.78 SubgraphRAG + GPT-4o-mini ($\leftrightarrow$) 74.42 49.41 89.10 84.67 81.35 49.47 45.16 60.18 54.18 58.86 SubgraphRAG + GPT-4o-mini ($\leftrightarrow$, 500) 76.83 52.01 92.30 87.43 81.41 55.58 49.13 67.24 59.69 59.87
4.2 KGQA Results (Q3 & Q4)
KGQA Baselines. Besides the baselines used for retriever evaluation, we include results from other LLM-based KGQA methods due to their state-of-the-art performance, such as KD-CoT (Wang et al., 2023a), StructGPT (Jiang et al., 2023a), ToG (Sun et al., 2024a), and EtD (Liu et al., 2024a). For RoG, we present two entries: RoG-Joint and RoG-Sep. Originally, RoG fine-tuned its LLMs on the training sets of both WebQSP and CWQ. Yet, this joint training approach leads to significant label leakage, with over 50% of WebQSP test questions (or their variants) appearing in CWQ’s training set, and vice versa. Therefore, we re-trained RoG on each dataset separately, indicated as RoG-Sep.
Evaluation Metrics. Along with the commonly reported Macro-F1 and Hit [^2] The baselines claim to report Hit@1, but they actually compute Hit, which measures whether at least one correct answer appears in the LLM response., we also include Micro-F1 to account for the imbalance in the number of ground-truth answers across samples and Hit@1 for a more inclusive evaluation. To further assess model performance, we introduce scoreh, inspired by (Yang et al., 2024), which evaluates how truth-grounded the predicted answers are and the degree of hallucination. This metric penalizes hallucinated answers while favoring missing answers over incorrect ones. Scores are normalized to a range of 0 to 100, with higher scores indicating better truth-grounding in the model’s answers. Details of the scoring strategy are provided in Appendix C.
Experiment Settings. We use the vllm (Kwon et al., 2023) framework for efficient LLM inference. For KD-CoT, Retrieve-Rewrite-Answer, ToG, and EtD, we report their published results due to the difficulty in reproducing them. For StructGPT, we directly use their provided processed files to obtain results for WebQSP. For the remaining baselines, we successfully reproduced their results and evaluated them on additional metrics and datasets, including WebQSP-sub and CWQ-sub. However, we do not include results for G-Retriever on CWQ, as it required over 200 hours of computation on 2 NVIDIA RTX 6000 Ada GPUs. If not specified, the LLM reasoners in SubgraphRAG use the top 100 retrieved triples; results using more triples are explicitly noted. All Llama variants considered are based on instruction tuning. For the LLM reasoners used in SubgraphRAG, both the temperature and the seed are set to 0 to ensure reproducibility.
@@ 4.2 KGQA Results (Q3 & Q4)
WebQSP-sub CWQ-sub $1$ $2$ $1$ $2$ $\geq 3$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit G-Retriever 56.41 78.20 45.73 65.35 - - - - - - RoG-Joint 77.05 92.96 62.53 81.54 59.75 66.33 59.70 68.56 41.46 43.27 RoG-Sep 74.50 89.83 55.62 73.45 59.35 66.20 59.45 67.17 31.33 33.33 SubgraphRAG (Llama3.1-8B) 75.50 91.40 65.87 83.62 51.54 63.05 57.52 68.93 41.88 47.37 SubgraphRAG (GPT-4o-mini) 80.56 92.86 74.11 88.51 57.36 67.34 63.85 72.74 51.14 54.39
@@
WebQSP CWQ WebQSP-sub CWQ-sub Macro-F1 Hit Macro-F1 Hit Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh SubgraphRAG + Rand 37.69 60.14 27.34 35.85 37.79 17.79 60.74 54.97 65.06 29.97 29.15 39.15 35.11 52.45 SubgraphRAG + RandNoAns 21.18 33.54 16.40 22.71 20.61 8.64 33.03 27.33 47.29 16.47 16.44 23.00 19.42 43.79 SubgraphRAG + NoRetriever 35.86 51.90 25.64 32.34 35.03 17.01 51.38 47.59 55.57 27.42 22.66 33.95 30.65 44.87 SubgraphRAG + cosine similarity 58.41 74.14 34.59 43.61 59.26 37.31 75.43 71.20 73.16 39.05 35.48 49.02 43.68 55.72 SubgraphRAG + Retrieve-Rewrite-Answer 8.96 11.43 - - 9.11 5.23 11.43 10.84 63.45 - - - - - SubgraphRAG + StructGPT 62.14 75.00 - - 62.55 44.72 75.69 73.57 80.82 - - - - - SubgraphRAG + G-Retriever 48.91 64.50 28.47 34.58 49.92 28.08 65.88 62.60 72.54 31.55 32.22 38.17 35.22 58.22 SubgraphRAG + RoG-Sep 57.68 74.39 36.85 45.23 59.36 40.32 76.65 72.61 79.69 44.11 43.69 54.04 47.68 66.18 SubgraphRAG 70.57 86.61 47.16 56.98 72.10 46.56 88.58 84.80 82.42 54.76 51.76 65.80 59.69 62.89
@@
Overall Performance. Tables 3 and 4 present the evaluation results, where SubgraphRAG achieves state-of-the-art (SOTA) results on both WebQSP and WebQSP-sub. Even with smaller 8B LLMs, our method surpasses previous SOTA approaches by up to 4% in Macro-F1 and Hit metrics (excluding RoG-Joint due to test label leakage in that model). With larger models like Llama3.1-70B-Instruct and GPT-4o, SubgraphRAG achieves even greater performance, showing up to a 12% improvement in Macro-F1 and a 9% increase in Hit. On the more challenging CWQ and CWQ-sub datasets, which require extended reasoning hops, SubgraphRAG performs competitively even with smaller 8B models. When paired with advanced reasoning models like GPT-4o, SubgraphRAG achieves results second only to ToG on CWQ. Notably, SubgraphRAG requires only a single call to GPT-4o, whereas ToG requires 6-8 calls, which increases computational cost, and we were unable to reproduce ToG’s performance using their published code. On CWQ-sub, SubgraphRAG demonstrates gains of up to 9% in Macro-F1 and 11% in Hit, indicating that tasks with greater reasoning complexity benefit substantially from more powerful LLMs. Our method also excels in truth-grounded QA, with Scoreh consistently outperforming baselines by up to 12%, driven by our prompt design that encourages explicit reasoning based on retrieved triples.
@@
In terms of efficiency, studies indicate that overall latency for an LLM to answer a question is primarily dominated by the number of LLM calls, followed by output token count, with input token count having minimal impact  [^3] For LLMs like GPT-4 and Claude-3.5, an additional input token typically adds about $10^{-2}\times$ the latency of an extra output token and $10^{-4}\times$ that of an additional LLM call (Vivek, 2024).. SubgraphRAG, despite potentially varying input token counts, utilizes only one LLM call per question, whereas various baselines require multiple calls.
Multi-Hop Performance Breakdown. Table 5 presents the performance breakdown by reasoning hops, with RoG as the primary baseline. Even with an 8B LLM, our method significantly outperforms RoG on multi-hop reasoning questions. This improvement can be attributed to our design approach, which avoids constraints on the types of retrieved subgraphs, allowing the LLMs’ reasoning capabilities to be better utilized. For 1-hop questions, SubgraphRAG outperforms the baselines on WebQSP, though its performance is slightly lower on CWQ.
Explainability. By retrieving flexible, high-quality evidence subgraphs as context and exploiting the strong reasoning capabilities of pre-trained LLMs, SubgraphRAG can natively provide effective explanations along with answer predictions. In contrast, explainable predictions with fine-tuned LLMs necessitate extra labeling efforts for preserving explainability (Luo et al., 2024) or post hoc explainability approaches (Krishna et al., 2023). To illustrate the explanability of our framework, we provide multiple examples of our reasoner’s responses (see Appendix E).
Performance with Different # of Retrieved Triples. It has been observed that irrelevant context can mislead LLMs (Wu et al., 2024), suggesting that retrieving more triples may not always improve performance. To investigate this, we conducted experiments (Fig. 4) using Llama3.1-8B-Instruct and GPT-4o-mini with varying numbers of retrieved triples. The upper bound is defined such that if the answer entity is present in the retrieved results, it is considered a correct answer. The results show that for the less powerful LLM, Llama3.1-8B-Instruct, performance initially improves with more triples but then declines, indicating the model is misled by irrelevant information and struggles to extract key triples from larger contexts. In contrast, GPT-4o-mini’s performance consistently improves with more triples, demonstrating greater robustness against irrelevant context. These findings suggest that LLMs have varying capacities for handling retrieved information, and the number of retrieved triples should be adjusted to align with the model’s capacity and available computational resources.
@@
Performance with Different Retrievers. To assess the impact of our retriever on the overall performance, we conducted extensive studies by replacing the retrieval results with those from baseline retrievers, while keeping all other settings constant (same prompts, same LLM reasoners). Tables 7 and 9 show the results using Llama3.1-8B-Instruct and GPT-4o-mini as the LLM reasoners, respectively. For retrievers from Retrieve-Rewrite-Answer, StructGPT, G-Retriever, and RoG, we used all their retrieved triples, while for other cases, we kept 100 triples. Clearly, pairing with our retriever yields significantly better performance on all metrics and datasets compared to using baseline retrievers, indicating that our superior results are largely due to the effectiveness of our improved retriever.
@@ 5 Conclusion
This paper presents SubgraphRAG, a novel framework for KG-based RAG. It performs efficient and flexible subgraph retrieval followed by prompting unfine-tuned LLMs for reasoning. SubgraphRAG demonstrates better or comparable accuracy, efficiency, and explainability compared to existing KG-based RAG approaches.
Acknowledgement
@@
Appendix A Additional Related Work
@@
With the rapid advancement of LLMs in recent years, researchers began to leverage them in KGQA. For instance, recent work has explored using LLMs as translators, converting natural language questions into executable SQL queries for KG databases to retrieve the answers (Jiang et al., 2023a; Wang et al., 2023b).
Contemporary approaches have further expanded the role of LLMs, utilizing them for both knowledge retrieval from KGs and reasoning (Kim et al., ; Gao et al., 2024a; Wang et al., 2024; Guo et al., 2024; Ma et al., 2024; Sun et al., 2024a; Jiang et al., 2024; Jin et al., 2024). The strength of this strategy lies in LLMs’ ability to handle multi-hop tasks by breaking them down into manageable steps. However, as discussed in Section 1, this often necessitates multiple LLM calls, resulting in high latency. To mitigate this issue, some frameworks have attempted to fine-tune LLMs to memorize knowledge, but this reduces their ability to generalize to dynamically updated or novel KGs (Luo et al., 2024; Mavromatis & Karypis, 2024). Other models have explored fine-tuning adapters embedded in fixed LLMs to better preserve their general reasoning capabilities while adapting to specific KGs (He et al., 2024; Gao et al., 2024a; Hu et al., 2024).
@@
Appendix B Additional Experiment Details for Subgraph Retrieval
B.1 Additional Implementation and Evaluation Details
For Retrieve-Rewrite-Answer, RoG, and G-Retriever, we utilize their official open-source implementations for training and evaluation. While a pre-trained RoG model is publicly available, it was jointly trained on the training subset of both WebQSP and CWQ, causing a label leakage issue due to sample duplication across the two datasets. To address this, we retrain the RoG model separately on the training subset of WebQSP and CWQ. Both Retrieve-Rewrite-Answer and G-Retriever were originally only evaluated on the WebQSP dataset. We managed to adapt the G-Retriever codebase for an evaluation on the CWQ dataset.
RoG, G-Retriever, the cosine similarity baseline, and all SubgraphRAG variants perform relevant subgraph retrieval from the identical rough subgraphs. These subgraphs are centered around the topic entities within a maximum number of hops, and they are included in the released RoG implementation. In contrast, Retrieve-Rewrite-Answer directly loads and queries the raw KG with a database server.

### v3

@@ 4.2 KGQA Results (Q3 & Q4)
WebQSP-sub CWQ-sub Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh SR+NSM w E2E 58.79 37.04 68.63 60.62 64.44 - - - - - G-Retriever 54.13 23.84 74.52 67.56 67.97 - - - - - RoG-Joint 72.01 47.70 88.90 82.62 76.13 58.61 52.12 66.22 61.17 55.15 RoG-Sep 67.94 43.10 84.03 77.61 72.79 57.69 52.83 64.64 60.64 54.51 SubgraphRAG + Llama3.1-8B 72.10 46.56 88.58 84.80 82.42 54.76 51.76 65.80 59.69 62.89 SubgraphRAG + Llama3.1-70B 75.97 51.64 87.88 85.89 85.57 61.49 59.91 68.43 65.52 67.62 SubgraphRAG + ChatGPT 70.81 44.73 85.18 80.82 81.53 56.37 54.44 64.40 60.99 61.31 SubgraphRAG + GPT4o-mini 78.34 58.44 91.34 87.36 82.21 61.13 58.86 70.01 65.48 64.20 SubgraphRAG + GPT4o 77.61 56.78 91.40 86.40 81.85 65.99 63.18 73.91 68.89 66.57 SubgraphRAG + GPT4o-mini (200) 78.66 58.65 91.73 87.04 81.98 61.58 57.47 71.45 65.87 63.66 SubgraphRAG + GPT4o (200) 79.40 58.91 92.43 87.75 82.46 66.48 61.30 75.14 69.42 66.45 SubgraphRAG + GPT4o-mini (500) 78.46 57.08 92.43 88.01 81.95 62.18 56.86 72.82 66.57 62.77 SubgraphRAG + Llama3.1-8B ($\leftrightarrow$) 67.91 42.79 85.25 81.21 80.09 43.03 40.73 55.09 47.58 56.78 SubgraphRAG + GPT4o-mini ($\leftrightarrow$) 74.42 49.41 89.10 84.67 81.35 49.47 45.16 60.18 54.18 58.86 SubgraphRAG + GPT4o-mini ($\leftrightarrow$, 500) 76.83 52.01 92.30 87.43 81.41 55.58 49.13 67.24 59.69 59.87 SubgraphRAG + GPT4o ($\leftrightarrow$) 74.98 50.67 88.26 84.73 83.24 52.40 52.28 60.36 56.39 63.70 SubgraphRAG + GPT4o ($\leftrightarrow$, 500) 77.96 56.40 90.96 86.66 83.35 61.96 58.48 70.47 64.96 66.86
@@
WebQSP-sub CWQ-sub $1$ $2$ $1$ $2$ $\geq 3$ $(65.8\%)$ $(34.2\%)$ $(28.0\%)$ $(65.9\%)$ $(6.1\%)$ Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit Marco-F1 Hit G-Retriever 56.41 78.20 45.73 65.35 - - - - - - RoG-Joint 77.05 92.96 62.53 81.54 59.75 66.33 59.70 68.56 41.46 43.27 RoG-Sep 74.50 89.83 55.62 73.45 59.35 66.20 59.45 67.17 31.33 33.33 SubgraphRAG (Llama3.1-8B) 75.50 91.40 65.87 83.62 51.54 63.05 57.52 68.93 41.88 47.37 SubgraphRAG (GPT4o-mini) 80.56 92.86 74.11 88.51 57.36 67.34 63.85 72.74 51.14 54.39
@@
WebQSP CWQ WebQSP-sub CWQ-sub Macro-F1 Hit Macro-F1 Hit Macro-F1 Micro-F1 Hit Hit@1 Scoreh Macro-F1 Micro-F1 Hit Hit@1 Scoreh Rand + SubgraphRAG Reasoner 37.69 60.14 27.34 35.85 37.79 17.79 60.74 54.97 65.06 29.97 29.15 39.15 35.11 52.45 RandNoAns + SubgraphRAG Reasoner 21.18 33.54 16.40 22.71 20.61 8.64 33.03 27.33 47.29 16.47 16.44 23.00 19.42 43.79 NoRetriever + SubgraphRAG Reasoner 35.86 51.90 25.64 32.34 35.03 17.01 51.38 47.59 55.57 27.42 22.66 33.95 30.65 44.87 cosine similarity + SubgraphRAG Reasoner 58.41 74.14 34.59 43.61 59.26 37.31 75.43 71.20 73.16 39.05 35.48 49.02 43.68 55.72 Retrieve-Rewrite-Answer + SubgraphRAG Reasoner 8.96 11.43 - - 9.11 5.23 11.43 10.84 63.45 - - - - - StructGPT + SubgraphRAG Reasoner 62.14 75.00 - - 62.55 44.72 75.69 73.57 80.82 - - - - - G-Retriever + SubgraphRAG Reasoner 48.91 64.50 28.47 34.58 49.92 28.08 65.88 62.60 72.54 31.55 32.22 38.17 35.22 58.22 RoG-Sep + SubgraphRAG Reasoner 57.68 74.39 36.85 45.23 59.36 40.32 76.65 72.61 79.69 44.11 43.69 54.04 47.68 66.18 GNN-RAG + SubgraphRAG Reasoner 62.34 78.81 46.40 54.35 63.93 42.43 80.89 77.10 82.86 55.08 53.38 64.33 59.76 67.55 GNN-RAG ($\leftrightarrow$) + SubgraphRAG Reasoner 52.07 71.38 34.88 43.13 53.01 34.58 72.93 67.99 73.37 39.98 34.60 49.54 44.91 57.42 SubgraphRAG 70.57 86.61 47.16 56.98 72.10 46.56 88.58 84.80 82.42 54.76 51.76 65.80 59.69 62.89 SubgraphRAG ($\leftrightarrow$) 66.42 83.42 37.96 48.57 67.91 42.79 85.25 81.21 80.09 43.03 40.73 55.09 47.58 56.78
@@
Experiment Settings. We use the vllm (Kwon et al., 2023) framework for efficient LLM inference. For UniKGQA, KD-CoT, Retrieve-Rewrite-Answer, ToG, and EtD, we reference their published results due to various reproducibility difficulties. For StructGPT, we directly use their provided processed files to obtain results for WebQSP. For SR, we include their reported results for WebQSP and CWQ, and further evaluate it on WebQSP-sub as we could not run through SR’s code on the two CWQ datasets. For the remaining baselines, we successfully reproduced their results and evaluated them on additional metrics and datasets, including WebQSP-sub and CWQ-sub. However, we do not include results for G-Retriever on CWQ as it required over 200 hours of computation on 2 NVIDIA RTX 6000 Ada GPUs. If not specified, the LLM reasoners in SubgraphRAG use the top 100 retrieved triples; results using more triples are explicitly noted. All Llama models used are their instruction fine-tuned versions, i.e., Llama3.1-8B(70B)-Instruct. The used GPT models’ versions are gpt-3.5-turbo-1106, gpt-4o-mini-2024-07-18, and gpt-4o-2024-08-06. For all LLM reasoners used in SubgraphRAG, both the temperature and the seed are set to 0 to ensure reproducibility.
Overall Performance. Tables 3 and 4 present the results, where SubgraphRAG achieves SOTA on WebQSP and WebQSP-sub. Even with 8B LLMs, our method surpasses previous SOTA approaches by $\leq 4\%$ in Macro-F1 and Hit metrics (excluding RoG-Joint due to test label leakage). With larger models like Llama3.1-70B-Instruct and GPT-4o, SubgraphRAG achieves greater performance, showing $\leq 12\%$ improvement in Macro-F1 and $\leq 9\%$ in Hit. On the more challenging CWQ and CWQ-sub datasets, which require more reasoning hops, SubgraphRAG performs competitively with smaller 8B models. When paired with advanced LLMs like GPT-4o, SubgraphRAG achieves results second only to ToG on CWQ. Notably, SubgraphRAG requires only a single call to GPT-4o, whereas ToG requires 6-8 calls, which increases computational cost, and we were unable to reproduce ToG’s performance using their published code. On CWQ-sub, SubgraphRAG demonstrates gains of $\leq 9\%$ in Macro-F1 and $\leq 11\%$ in Hit, indicating that tasks with greater reasoning complexity benefit substantially from more powerful LLMs. Our method also excels in truth-grounded QA, with Scoreh consistently outperforming baselines by $\leq 12\%$, driven by our prompt design that encourages retrieval-grounded reasoning. See Appendix D for further discussions.
@@
For efficiency, studies show that overall LLM latency is dominated by ##\## LLM calls, followed by ##\## output tokens and then ##\## input tokens  [^2] For LLMs like GPT-4 and Claude-3.5, an additional input token typically adds about $10^{-2}\times$ the latency of an extra output token and $10^{-4}\times$ that of an additional LLM call (Vivek, 2024).. SubgraphRAG, despite potentially varying ##\## input tokens, utilizes only one LLM call per question, whereas various baselines require multiple calls.
Performance Breakdown & Analysis. Table 5 presents the performance breakdown by reasoning hops. Our method, even with an 8B LLM, shows significantly better results on multi-hop questions. This can be attributed to our design philosophy, which does not restrict the retrieved subgraph form and leverages unfine-tuned LLMs with ICL examples, fully unlocking their reasoning capabilities. Interestingly, SubgraphRAG shows relatively lower performance on some 1-hop questions in CWQ-sub. Cross-referencing with Table 8, which provides a detailed breakdown of whether the answers are truth-grounded, we find that previous methods may generate many correct answers that are not supported by the retrieved results. Further investigation suggests this discrepancy may partly result from dataset leakage, giving tuning-based methods an advantage: In CWQ, over 60% of test samples have answers that have appeared in the training set, making it easier for these methods to succeed without robust reasoning. Thus, future studies are needed to fully investigate the potential leakage.
@@
Performance with Different Retrievers. To assess the impact of our retriever on the overall performance, we conducted extensive studies by replacing the retrieval results with those from baseline retrievers, while keeping all other settings constant (same prompts, same LLM reasoners). Tables 6 and 9 show the results using Llama3.1 8B and GPT4o-mini as the LLM reasoners, respectively. For retrievers from Retrieve-Rewrite-Answer, StructGPT, G-Retriever, and RoG, we used all their retrieved triples, while for other cases, we kept 100 triples. Clearly, pairing with our retriever yields significantly better performance on all metrics and datasets compared to using baseline retrievers, indicating that our superior results are largely due to the effectiveness of our improved retriever.
Impact of Retrieval Size. Irrelevant context can mislead LLMs (Wu et al., 2024), and retrieving more triples may not improve performance. To investigate this, we conducted experiments (Fig. 4) by feeding Llama 3.1 8B and GPT-4o-mini different numbers of retrieved triples. The results show a significant performance decay in Llama 3.1 8B with more triples, while GPT-4o-mini performs even better with additional retrieved triples. This suggests that different LLMs have varying capacities for robust reasoning, and users should input proper numbers of retrieved results to match the capacity of the used LLM and the available computational budget. See Appendix D for more results.
Explainability. By retrieving flexible, high-quality evidence subgraphs as context and exploiting the strong reasoning capabilities of pre-trained LLMs, SubgraphRAG can natively provide effective explanations along with answer predictions. In contrast, explainable predictions with fine-tuned LLMs necessitate extra labeling efforts for preserving explainability (Luo et al., 2024b) or post hoc explainability approaches (Krishna et al., 2023). To illustrate the explanability of our framework, we provide multiple examples of our reasoner’s responses (see Appendix F).
@@ 5 Conclusion
We propose SubgraphRAG, a KG-based RAG framework that performs efficient and flexible subgraph retrieval followed by prompting unfine-tuned LLMs. SubgraphRAG demonstrates better or comparable accuracy, efficiency, and explainability compared to existing approaches.
Acknowledgments
@@
Appendix A Directional Distance Encoding (DDE)
Figure 5: A visual illustration of DDE, where the leftmost column highlights the topic entities in color. From left to right, we iteratively perform graph diffusion over the directed edges. The color intensity indicates the magnitude of the diffusion weights for individual entities.
Fig. 5 provides a visual illustration of DDE. By passing entity labeling information in both directions for multiple rounds, DDE efficiently approximates measures such as the average distance between an entity and all topic entities. This process implicitly reveals information like whether an entity lies at the intersection of paths of particular lengths connecting multiple topic entities.
Appendix B Additional Related Work
@@
With the rapid advancement of LLMs in recent years, researchers began to leverage them in KGQA. For instance, recent work has explored using LLMs as translators, converting natural language questions into executable SPARQL queries for KG databases to retrieve the answers (Jiang et al., 2023a; Wang et al., 2023b; Luo et al., 2024a; Xiong et al., 2024).
Contemporary approaches have further expanded the role of LLMs, utilizing them for both knowledge retrieval from KGs and reasoning (Kim et al., ; Gao et al., 2024a; Wang et al., 2024; Guo et al., 2024; Ma et al., 2024; Sun et al., 2024a; Jiang et al., 2024; Jin et al., 2024). The strength of this strategy lies in LLMs’ ability to handle multi-hop tasks by breaking them down into manageable steps. However, as discussed in Section 1, this often necessitates multiple LLM calls, resulting in high latency. To mitigate this issue, some frameworks have attempted to fine-tune LLMs to memorize knowledge, but this reduces their ability to generalize to dynamically updated or novel KGs (Luo et al., 2024b; Mavromatis & Karypis, 2024). Other models have explored fine-tuning adapters embedded in fixed LLMs to better preserve their general reasoning capabilities while adapting to specific KGs (He et al., 2024; Gao et al., 2024a; Hu et al., 2024).
@@
Appendix C Additional Experiment Details for Subgraph Retrieval
C.1 Additional Implementation and Evaluation Details
For Retrieve-Rewrite-Answer, RoG, G-Retriever, GNN-RAG, and SR+NSM w E2E, we utilize their official open-source implementations for training and evaluation. While a pre-trained RoG model is publicly available, it was jointly trained on the training subset of both WebQSP and CWQ, causing a label leakage issue due to sample duplication across the two datasets. To address this, we retrain the RoG model separately on the training subset of WebQSP and CWQ. Both Retrieve-Rewrite-Answer and G-Retriever were originally only evaluated on the WebQSP dataset. We managed to adapt the G-Retriever codebase for an evaluation on the CWQ dataset. While SR+NSM w E2E originally also reports the results for CWQ, we failed in running the official codebase for this dataset.
RoG, G-Retriever, the cosine similarity baseline, GNN-RAG, and all SubgraphRAG variants perform relevant subgraph retrieval from the identical rough subgraphs. These subgraphs are centered around the topic entities and are included in the released RoG implementation. In contrast, Retrieve-Rewrite-Answer directly loads and queries the raw KG.